The demand for internet services in the town of Knoxville is given by Q = 10000 − 50P, where Q is the number of households serviced and P is the price of the service per month. The marginal cost of providing internet services per household is $10.

a. If Verizon is the only provider of internet services in the town, what price can it set and how many households would be served?

b. Now suppose a new firm INSAT enters the market. What happens to the price and output? How is the profit of verizon affected?
Answer #1

Q = 1000D - 5Up => 50 P = 19000Q = P = 20 -0.020 mc = 10 TC=100 6) TR = P Q = 20Q -0.020² MR = 20-0.040 het mR=mc to maximize